# Stark localization near Aubry-André criticality

### v1

@@ IV Scaling Properties near AA criticality
In the previous section, we have done the scaling analysis at the AA criticality, i.e., $\delta=0$. Now we turn our attention to the near criticality region of the AA model. Now as $\delta\neq 0$ one can expect that the previous scaling ansatz will not work properly and some modifications are needed.
@@
Figure 5: Ground State Wavefunction $\psi_{0}$: (a1) describes the ground state wavefunction for the pure AA model, where $h=0$ and $\delta=0.4J$. (a2) presents the ground state wavefunction for the pure stark model, where $\delta=-2J$ and $h=10^{-4}$. (a3) presents the ground state wavefunction for $\delta=0$ with $h=10^{-4}$ and (a4) presents the wave function for $\delta=-0.1$ and $h=10^{-4}$. All plots are done for system size 610 and for the AA potential, the phase, $\phi=0$.
To examine the validation of the above scaling ansatz, we now turn our attention to delve into scaling properties of $\zeta$, IPR, and $\Delta E$ near criticality for $\delta L^{1/\nu}=1$ and $\delta L^{1/\nu}=-1$. Upon taking $\delta$ in this fashion for different $L$, we expect that our previous critical exponents would work as we are still in the vicinity of the AA critical point. In Fig. 2 both these cases are shown and we have numerically found that the collapse still happens for the same set of scaling exponents as found in the case of $\delta=0$.
V Scaling Analysis at Far from the Aubry-André Criticality
@@
In Fig. 3(a1), the localization length ($\zeta$) is plotted against $h$, while in Fig. 3(a2), we depict $\zeta/L$ against $hL^{1/\nu}$, with a specific value of $\nu=0.33$. Remarkably, all datasets collapse onto a single curve for this value of $\nu$. This observation suggests that in the as we move away from the AA criticality the scaling nature replicates the scaling of Stark localization. The IPR for $\delta=-0.1$ across various system sizes is shown in Fig. 3(b1). Scaling analysis for IPR gives the value, $s=0.33$, which is confirmed by data collapse shown in Fig. 3(b2) and also by the cost-function approach. This scaling exponent is identical to that of the pure Stark model. In Fig. 3(c1), we plot the dependence of $\Delta E$ on $h$. The nature of the dependence is found to be the same as before. The collapse plot corresponding to $\Delta E$ is depicted in Fig. 3(c2), yielding a critical exponent of $z=2.00(2)$ which is again the same critical exponent value for that of the pure Stark model.
@@
Thus, as we move away from the AA criticality the system does not experience much effect due to the AA potential. Likewise, the delocalization-localization transition is solely controlled by the Stark field. Therefore we study the behaviour of the scaling exponents as we slowly move away from the criticality. Fig. 4 nicely depicts this variation. The variation of $s/\nu$ and $z$ are shown in Fig. 4(a1) and Fig. 4(a2) respectively. To be specific, here, we only focus on the negative values of $\delta$, as we have found that a slight departure from the critical point in the positive $\delta$ values results in a very weird kind of variation in the characteristic observables and as a consequence of that, the scaling analysis fails there. This happens due to the competition between the two kinds of localization. From Fig. 4 it is easily noticeable that in the negative $\delta$ region, there is a continuous change in the values of the scaling exponents. As we move away from the AA criticality, the values of $s/\nu$ and $z$ slowly change from that of the pure AA model to the pure Stark model.
Moreover, in order to understand the competing nature of two distinct types of localization inducing potentials closely, it is interesting to directly look into the structures of the ground state wavefunctions for various cases. We present some exemplary cases in Fig. 5 for various scenarios relevant to our study of the scaling nature of the system near the criticality.
@@ VI Hybrid Scaling Exponent
VII CONCLUSION
@@
In this work we have explored the Stark localization in the presence of a quasi-periodic potential, namely the AA potential. This exploration is based on the scaling analysis of finite-size systems. First, we have studied the effect of increasing Stark field at the AA criticality. We have found that the scaling exponents $\nu$ and $s$ are quite different than both the pure AA model and the pure Stark model. However, the $s/\nu$ value is found to be the same as that of the pure AA model. Also, the scaling exponent $z$ is identical to that of the pure AA model. It indicates although the Stark field drives the transition, the transition is mostly AA type. Next, we have studied the same near the AA criticality. As in that case, we have used $\delta\neq 0$, we cannot employ the scaling ansatz for one control parameter-based phase transition. We have developed some scaling ansatz that is capable of extracting the relevant scaling exponents quite efficiently. Then, we take a fixed system size and vary the distance from the AA criticality. For this, a hybrid scaling ansatz is used and it successfully describes the scaling nature. Finally, we have looked into the variation of $s/\nu$ and $z$ with negative values of $\delta$ which have revealed the fact that near the AA criticality, these exponents coincide with that of the pure AA model, and as we move away from the AA criticality we enter into the region where the localization is solely driven by Stark field with almost no trace of the presence of AA potential.

### v2

@@ IV Scaling Properties near AA criticality
In the previous section, we have done the scaling analysis at the AA criticality, i.e., $\delta=0$. Now we turn our attention to the near criticality region of the AA model. Now as $\delta\neq 0$, one can expect that the previous scaling ansatz will not work properly and some modifications are needed.
@@
To examine the validation of the above scaling ansatz, we now turn our attention to delve into scaling properties of $\zeta$, IPR, and $\Delta E$ near criticality for $\delta L^{1/\nu}=1$ and $\delta L^{1/\nu}=-1$. Upon taking $\delta$ in this fashion for different $L$, we expect that our previous critical exponents would work as we are still in the vicinity of the AA critical point. In Fig. 3 both these cases are shown and we have numerically found that the collapse still happens for the same set of scaling exponents as found in the case of $\delta=0$. The value of $\nu$ obtained from fitting for $\delta L^{1/\nu}=1$ is $0.295(4)$. For $\delta L^{1/\nu}=-1$, the value of $\nu$ is $0.297(2)$. For $\delta L^{1/\nu}=1$, the value of $s$ is $0.095(2)$, while for $\delta L^{1/\nu}=-1$, the value of $s$ is $0.099(2)$. Based on the cost function approach for $\delta L^{1/\nu}=1$, the value of $\nu$ lies in the range $0.285$ to $0.298$, giving an average of $0.292(1)$. The corresponding value of $s$ obtained is $s=0.0964(1)$. Similarly, for $\delta L^{1/\nu}=-1$, the value of $\nu$ is in the range $0.285$ to $0.298$, with an average of $0.291(5)$, and the value of $s$ obtained is $s=0.0972(1)$.
Figure 5: (a1) presents $s/\nu$ for different values of $\delta$. For $\delta=0$ the $s/\nu$ value is $0.33$ which is the same as that of the pure AA model and this value continuously increases to $1$, which is the same as the pure Stark model, as one moves away from the AA criticality with negative $\delta$ values. (a2) presents the plot of $z$ for different values of $\delta$. In this case, there is a continuous change in the value of $s/\nu$ from 2.37, which is the same as the pure AA model, to 2, which is for the pure Stark model, as one moves away from the criticality.
Figure 6: Ground State Wavefunction $\bm{\psi_{0}:}$ (a1) describes the ground state wavefunction for the pure AA model, where $h=0$ and $\delta=0.4J$. (a2) presents the ground state wavefunction for the pure stark model, where $\delta=-2J$ and $h=10^{-4}$. (a3) presents the ground state wavefunction for $\delta=0$ with $h=10^{-4}$ and (a4) presents the wave function for $\delta=-0.1$ and $h=10^{-4}$. All plots are done for system size 610 and for the AA potential, the phase, $\phi=0$.
Figure 7: We show the fidelity $\mathcal{F}(h,\delta)$ between the ground state of the AAS system (i.e., the system under the dual potential), say $|\psi_{0}(h,\delta)\rangle$, with the ground state of the pure stark model, i.e., $|\psi_{0}(h,\delta=0)\rangle$, in the parametric space of $h$ and $\delta$ for system size $L=610$. The range of the quantitative value is shown in the color bar. We have performed a configuration averaging over $\phi$ with 100 random samples. The blue region corresponds to the parametric space with high fidelity (90% or more) indicating negligible influence of the AA potential, and hence, implies Stark dominated localization.
V Scaling Analysis at Far from the Aubry-André Criticality and Stark dominated localization
@@
In Fig. 4(a1), the localization length ($\zeta$) is plotted against $h$, while in Fig. 4(a2), we depict $\zeta/L$ against $hL^{1/\nu}$, with a specific value of $\nu=0.33$. Remarkably, all datasets collapse onto a single curve for this value of $\nu$. From the fitting function in 4(a1), $\zeta\sim h^{-\nu}$, We get $\nu=0.337(5)$. This observation suggests that as we move away from the AA criticality the scaling nature replicates the scaling of Stark localization. The IPR for $\delta=-0.1$ across various system sizes is shown in Fig. 4(b1). Scaling analysis for IPR gives the value, $s=0.347(3)$, which is confirmed by data collapse shown in Fig. 4(b2) and also by the cost-function approach. This scaling exponent is identical to that of the pure Stark model. In Fig. 4(c1), we plot the dependence of $\Delta E$ on $h$. The nature of the dependence is found to be the same as before. The collapse plot corresponding to $\Delta E$ is depicted in Fig. 4(c2), yielding a critical exponent of $z=2.00(2)$ which is again the same critical exponent value for that of the pure Stark model. In Fig. 4(c1) from the scaling of $\Delta E$ with $h$, the green color fitting function gives the value $\nu z=0.682(2)$ with $0.32\%$ error and the value of $z$ turns out to be $1.989(1)$. From the cost function, the range of $\nu$, for which $C_{Q}$ remains minimal and flat is from $0.326$ to $0.344$. The average of $\nu$ within this window is $0.335$. Finally, the extracted values of $s$ and $z$ are $0.03457(1)$ and $1.997(1)$, respectively.
@@
Thus, as we move away from the AA criticality the system does not experience much effect due to the AA potential. Likewise, the delocalization-localization transition is dominated by the Stark field. Therefore we study the behaviour of the scaling exponents as we slowly move away from the criticality. Fig. 5 nicely depicts this variation. The variation of $s/\nu$ and $z$ are shown in Fig. 5(a1) and Fig. 5(a2) respectively. To be specific, here, we only focus on the negative values of $\delta$, as a slight departure from the critical point in the positive $\delta$ values results in an unusual variation in the characteristic observables due to the finite localization of the underlying AAH potential, and as a consequence, the scaling analysis fails in this region. This happens due to the competition between the two kinds of localization. From Fig. 5 it is easily noticeable that in the negative $\delta$ region there is a continuous change in the values of the scaling exponents. As we move away from the AA criticality, the values of $s/\nu$ and $z$ slowly change from that of the pure AA model to the pure Stark model.
Moreover, in order to understand the competing nature of two distinct types of localization-inducing potentials closely, it is interesting to directly look into the structures of the ground-state wavefunctions for various cases. We present some exemplary cases in Fig. 6 for various scenarios relevant to our study of the scaling nature of the system near the criticality. The reflection symmetry of the system is strongly broken in the presence of the Stark potential, and, correspondingly, the ground-state localized wavefunction in the pure Stark of the system is strongly peaked near one of the ends of the lattice, unlike the pure AA case. Setting the AA potential near the AA criticality introduces exotic structures characterized by small peaks around a central large peak, because of the strong influence of the AA potential. As $|\delta|$ increases, i.e., for which the pure AA system is delocalized, and as $h$ increases, the effects of the AA potential in the AAS system gradually diminish, as one may expect. In order to explain this, which has also been presented in the schematic diagram Fig. 1, we compute the quantity, $\mathcal{F}(h,\delta)=|\langle\psi_{0}(h,\delta)|\psi_{0}(h)\rangle|$, which is the fidelity between the ground state of the system of AAS model $|\psi_{0}(h,\delta)\rangle$ and that of the pure Stark model $|\psi_{0}(h,\delta=0)\rangle$ and provides an estimate on the closeness of these two quantum states. The high value of $\mathcal{F}$ essentially implies that the underlying physics of the AAS model is effectively governed by the Stark potential. In Fig. 7, We have shown the bicolor plot, whose x axis is $-\delta$, the y axis is $h$ and the colored region is described by $\mathcal{F}(h,\delta)$. The blue region is where $\mathcal{F}(h,\delta)\geq 0.9$. The localization in this region is essentially stark dominated. The yellow region is for $0\leq\mathcal{F}(h,\delta)\leq 0.9$. The AA potential has a strong influence near $|\delta|\to 0$ and $h=0$, which gradually decreases with increasing values of $h$ and $\delta$. As discussed in the previous paragraph, in the limit of $h\to 0$, the manifestation of Stark dominated has been demonstrated by tracking the evolution of the scaling exponents with the change of $|\delta|$. In addition, the system is sufficiently away from the AA criticality ($|\delta|=0.1$), the scaling exponents turn out to be identical to the pure Stark system. As one may expect, $\mathcal{F}(h,\delta)$ is quite high for $|\delta|\geq 0.1(\mathcal{F}\gtrsim 0.84\%)$.
@@ VI Hybrid Scaling Exponent
VII Application as quantum sensor
Quantum sensing, a crucial pillar of quantum technology, enables the estimation of unknown parameters, e.g., magnetic fields, electric fields, gravitational fields, time, and other physical quantities, with remarkable precision. Recent experimental advancements towards developing quantum sensing devices, which include interferometry interferometry ; Nagata07 , magnetometry Wasilewski10 ; Sewell12 , and ultracold spectroscopy Liebfried04 ; Roos06 , among others, demonstrate the advantages originating from the quantum domain over the conventional classical counterparts. Recently, several works have proposed criticality-based adiabatic quantum sensing devices. The motivation behind leveraging criticality for developing quantum sensors lies in the fact that a many-body system is susceptible to small changes in a system parameter near criticality Zanardi08 . In the context of measuring an unknown field, it has been demonstrated that QFI, $F_{Q}$ (see below for a formal definition), can saturate the Heisenberg limit (HL) at the localization-delocalization transition in the pure AA model, implying a quantum advantage in precision measurement Sahoo24 . Similarly, localization transition for pure stark model enables $F_{Q}$ to beat HL He23 , and hence, paves the way towards designing an efficient quantum weak-field sensor. Here, our goal is to show that a better quantum sensor can be engineered by introducing an additional control field in the form of the AA potential for precision estimation of the Stark weak field. In the following, we briefly discuss the idea of quantum sensing and present the formal definition of the QFI, $F_{Q}$.
One needs to perform measurements for the estimation of an unknown parameter $h$. Let’s consider that the parameter $h$ is encoded in a quantum state $\rho(h)$. In the projective measurement basis $\{\pi_{n}$}, the probability corresponding to the $n^{\text{{th}}}$ outcome is given by, $p_{n}(h)=\text{Tr}[\rho(h)\pi_{n}]$. In this situation, the standard deviation of $h$ is bounded by the relation $\delta h\geq 1/\sqrt{MF_{c}}$ Cramer46 ; Helstrom76 , where $M$ is the number of measurements, and the Classical Fisher information, $F_{c}$ is defined as,
$F_{c}=\sum_{n}\frac{1}{p_{n}(h)}\left(\frac{dp_{n}(h)}{dh}\right)^{2},$ (17)
Figure 9: The scaling exponents $\nu$, $s$ and $z$ at and near the AAH criticality are tabulated.
Figure 10: Single party QFI: (a) presents QFI against the stark field $h$ for system size L = 144 (blue circle), 233 (red square), 377 (turquoise diamond), 610 (black triangular up), 987 (green triangular left). (b) shows the scaling plot. The maroon circles are QFI for different system sizes $L=21,34,55,89,144,233,377,610$ at $h=10^{-9}$, and the dotted line is the best fit $F_{Q}\sim L^{6.7}$. Both plots are shown in log-log scale. A configuration averaging over $\phi$ is performed with 8000 random samples.
Maximization over all possible measurements, one can obtain a measurement-independent quantity, the so-called QFI, $F_{Q}$. The QFI can be expressed as, $F_{Q}=\text{Tr}[L_{h}^{2}\rho(h)]$, where $L_{h}$ is the symmetric logarithmic derivative that is defined implicitly as $\partial_{h}\rho(h)=1/2\{\rho(h),L_{h}\}$. For a pure state, $\rho(h)=|\psi(h)\rangle\langle\psi(h)|$, the $L_{h}$ can be written as $2\partial_{h}\rho(h)$. The QFI now turns out to be,
$F_{Q}=4(\langle\partial_{h}\psi(h)|\partial_{h}\psi(h)\rangle-|\langle\partial_{h}\psi(h)|\psi(h)\rangle|^{2}).$ (18)
In the many-body context, system size is a resource, and hence, the scaling of QFI, $F_{Q}$, with system size, $L$, is of key interest Zanardi08 .
Considering QFI scales as, $F_{Q}\sim L^{\beta}$, $\beta=1$ corresponds to the standard quantum limit (SQL), which is what a classical system can achieve at best with $L$ probes. A definitive quantum advantage corresponds to a case, $\beta>1$. The special case of $\beta=2$ is the so-called Heisenberg scaling, beyond which, i.e., for $\beta>2$, is termed as super-Heisenberg scaling. Recently, it has been shown that the Stark transition can be exploited for precision measurement of an unknown weak field He23 . There the ground state QFI scales as $F_{Q}\sim L^{5.9}$, and hence, achieves a super-Heisenberg scaling with the system size. In the following, as mentioned before, we discuss whether it is possible to obtain an enhanced scaling of the QFI in the presence of an additional AAH potential that we consider as a control field, and thus it is possible to design a better quantum sensor for weak field estimation.
We consider the control field, i.e., the AA potential is set at AA criticality, i.e., $\delta=0$. Considering the ground state as the probe state, we compute QFI using Eq. (18). In Fig. 10, we present $F_{Q}$ as a function of the stark strength, $h$, for various system sizes $L=144,233,377,610,987$. Throughout the calculations, averaging over $\phi$ has been performed with OBC. The figure illustrates that, for finite systems, as the Stark strength $h$ gradually increases, the QFI remains flat in the extended region. However, beyond this flat region, the QFI decreases in the localized region. Evidently, in the flat areas of the finite systems, the QFI has a finite-size scaling, whereas it becomes scale-invariant in the localized phase, as one may expect. Fig. 10(b) illustrates the scaling of QFI with system size in the extended region. The maroon circles represent QFI values for different system sizes, and the maroon dashed line shows the fitting function $F_{Q}\sim L^{\beta}$, where $\beta$ is 6.70(1). Evidently, the $F_{Q}$ at the transition point scales as $F_{Q}\sim L^{\beta}$, where $\beta$ turns out to be $\beta=2/(d\nu)$ with $d$ as the spatial dimension. The scaling exponent $\nu\sim 0.29$ is associated with the localization length of the hybrid system under the influence of two different kinds of localizing potentials (see Sec. III). It’s worth mentioning here that, in the context of the second-order phase transition, the QFI is characterized by $\nu$ – the scaling exponent associated with the localization length. In the localization transition induced many-body quantum sensors, the scaling exponent associated with the localization length plays the same role in the QFI. The enhanced scaling of QFI in the presence of the AA criticality establishes the superiority of the system under consideration as a many-body quantum sensor for precision measurement of the weak field in comparison to the pure Stark sensors.
Figure 11: Cost function for $\delta=0$: (a1) presents plot for cost function with respect to $\nu$. It shows that the $C_{Q}$ has global minima for $\nu=0.29$. The inset shows an enlargement near the global minima. (a2) shows plot $C_{Q}$ versus $s$ and the minimum value of $C_{Q}$ is for $s=0.096$. (a3) demonstrates the plot $C_{Q}$ versus $z$. The minimum value of $C_{Q}$ is at $z=2.37$. Cost function for $\delta=-0.1$: (b1) presents $C_{Q}$ versus $\nu$. The $C_{Q}$ has the lowest value for $\nu=0.33$. (b2) shows the plot for $C_{Q}$ versus $s$. Here also at $s=0.33$, the $C_{Q}$ has global minima. (b3) displays the $C_{Q}$ versus $z$ and the minimum value of $C_{Q}$ is for $z=2$.
VIII CONCLUSION
@@
In this work we have explored the Stark localization in the presence of a quasi-periodic potential, namely the AA potential. This exploration is based on the scaling analysis of finite-size systems. First, we have studied the effect of increasing Stark field at the AA criticality. We have found that the scaling exponents $\nu$ and $s$ are quite different than both the pure AA model and the pure Stark model. The value of $s/\nu$, however, is found to be the same as that of the pure AA model. Also, the scaling exponent $z$ is identical to that of the pure AA model. It indicates although the Stark field drives the transition, the transition is mostly AA type. Next, we have studied the same near the AA criticality. As in that case, we have used $\delta\neq 0$, we cannot employ the scaling ansatz for one control parameter-based phase transition. We have developed some scaling ansatz that is capable of extracting the relevant scaling exponents quite efficiently. We tabulate the critical exponents at and near the AAH potential in Fig. 9 as a summary. In addition, we consider a fixed system size and vary the distance from the AA criticality. For this, a hybrid scaling ansatz is used and it successfully describes the scaling nature. It is evident from our numerical results that critical exponent violates Harris-Chayes criterion Harris74 ; Chayes86 which suggests $\nu\geq 2/d=2$, where $d$ is a spatial dimension. In this respect it may be interesting to investigate the scope of a field theoretical treatment, e.g., via real space renormalization group theory, in providing an alternate route for the understanding of the localization transition in such systems under study. Finally, we investigate the variation of $s/\nu$ and $z$ with negative values of $\delta$, which reveals the fact that near the AA criticality, these exponents coincide with that of the pure AA model, and as we move away from the AA criticality we enter into the region where the localization is solely driven by Stark field with almost no trace of the presence of AA potential. We complement these studies with a fidelity based approach in order to gauge the parametric regime, where the underlying localization physics is essentially Stark dominated. These studies find an immediate application towards engineering an efficient Stark weak field sensor. We have shown that the presence of an AA control field results in an enhanced scaling of the QFI in comparison to the pure Stark model, establishing the hybrid system as a better sensing device. A similar idea may be useful in other contexts, in general, e.g., using additional control parameters for designing better criticality-based sensing devices across various kinds of quantum phase transitions.
Appendix A Cost Function
The scaling exponents are extracted via a cost function, $C_{Q}$, approach, where $C_{Q}$ is defined as Suntajs20 ,
In the following, we present certain representative cases demonstrating the behavior of the cost function around the estimated critical exponents. We plot $C_{Q}$ with respect to the critical exponents $\nu$, $s$ and $z$ in Fig. 11. The Fig. 11(a1-a3) demonstrate the plot for $C_{Q}$ versus $\nu$, $s$ and $z$, respectively for $\delta=0$. The figures depict that the $C_{Q}$ has a global minimum and the minimum value of $C_{Q}$ is at $\nu=0.29$, $s=0.096$, and $z=2.37$, these values agree the value of scaling exponents that extract from collapse plot. Similarly for $\delta=-0.1$, the $C_{Q}$ versus $\nu$, $s$, and $z$ are shown in Fig. 11(b1-b3), respectively. Here also $C_{Q}$ has a local minimum. All data collapses for the minimum value of $C_{Q}$ at $\nu=0.33$, $s=0.33$, and $z=2$ for $\delta=-0.1$.